### v3

@@ Abstract.
In this paper, we introduce the $F$-invariant in cluster algebras using tropicalization. This is an analog of the $E$-invariant introduced by Derksen, Weyman and Zelevinsky in the additive categorification of cluster algebras and the $\mathfrak{d}$-invariant introduced by Kang, Kashiwara, Kim and Oh in the monoidal categorification of (quantum) cluster algebras. We prove that the product of two cluster monomials is still a cluster monomial if and only if their $F$-invariant is zero. For cluster algebras with a compatible Poisson bracket, we prove that if two cluster variables are log-cannonical, then they are contained in the same cluster.
Inspired by $F$-invariant, we introduce the dominant sets for seeds of cluster algebras as an replacement of torsion classes for $\tau$-tilting pairs in $\tau$-tilting theory and as an replacement of inversion sets in the study of right weak order on Weyl groups. With the help of the dominant sets, we prove that the oriented exchange graphs of cluster algebras are acyclic. In particular, this implies that green mutations induce a partial order on the set of seeds (up to seed equivalence) of cluster algebras. We prove that the oriented exchange graphs of cluster algebras coincide with the Hasse quivers of the above posets of seeds.
@@ 1. Introduction
Cluster algebras. Cluster algebras were invented by Fomin and Zelevinsky [fz_2002] as a combinatorial approach to the upper global bases of quantum groups and to the theory of total positivity in algebraic groups. Such algebras often arise as the coordinate rings of spaces, such as double Bruhat cells [bfz_2005], unipotent cells [gls_2011], double Bott-Samelson cells [Shen-Weng-2021] and arise as the Grothendieck rings of certain monoidal subcategories of the representations of quantum affine algebras [HL_2010, HL_2013].
@@
Compatibility degrees in cluster algebras. In the study of Y-systems and generalized associahedra, Fomin and Zelevinsky [fz-2003y] introduce a function
$(-\mid\mid-):\Phi_{\geq-1}\times\Phi_{\geq-1}\rightarrow\mathbb{Z}_{\geq 0}$
defined on the set $\Phi_{\geq-1}$ of almost positive roots associated to a Cartan matrix $C$ of finite type. Let $\mathcal{A}$ be a cluster algebra of finite type with the same Cartan-Killing type as the Cartan matrix $C$. Fomin and Zelevinsky [fz_2003f] proved that there is a bijection from $\Phi_{\geq-1}$ to the set $\mathcal{X}$ of cluster variables of $\mathcal{A}$ and they proved that two cluster variables $x_{\alpha}$ and $x_{\beta}$ are contained in the same cluster if and only if their corresponding roots $\alpha,\beta\in\Phi_{\geq-1}$ are compatible, i.e., $(\alpha\mid\mid\beta)=0$.
Now let $\mathcal{A}$ be any skew-symmetrizable cluster algebra (not necessarily finite type). In [cao-li-2020], Li and the author of the present paper proved that there exists a well-defined function
$(-\mid\mid-)_{d}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{Z}_{\geq-1}$
defined on the set of cluster variables $\mathcal{X}$ of $\mathcal{A}$, which is called the $d$-compatibility degree. The values of the $d$-compatibility degree $(-\mid\mid-)_{d}$ are given by the components of the $d$-vectors [fomin_zelevinsky_2007] in cluster algebras. It is proved that two cluster variables $x,z\in\mathcal{X}$ are contained in the same cluster if and only if
@@
$(-\mid\mid-)_{f}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{Z}_{\geq 0}$
@@
defined on the set $\mathcal{X}$ of cluster variables of $\mathcal{A}$. Based on the results in [cao-li-2020], Fu and Gyoda proved that two cluster variables $x,z\in\mathcal{X}$ are contained in the same cluster if and only if $(x\mid\mid z)_{f}=0$. We remark that when restricting to cluster algebras of finite type, the $f$-compatibility degree coincide with Fomin-Zelevinsky’s compatibility degree.
$E$-invariant. The $E$-invariant is introduced by Derksen, Weyman and Zelevinsky [DWZ10] in the additive categorification of cluster algebras using decorated representations of quivers with potentials. We refer to [DWZ08, Definition 10.1] for the definition of decorated representations. For any two decorated representations $\mathcal{M}$ and $\mathcal{N}$ of a quiver with potential $(Q,W)$, Derksen, Weyman and Zelevinsky [DWZ10, Section 7] defined an integer $E^{\rm inj}(\mathcal{M},\mathcal{N})$, which turns out to take non-negative values (see [DWZ10, Section 10]). The $E$-invariant $E^{\rm sym}(\mathcal{M},\mathcal{N})$ associated to $\mathcal{M}$ and $\mathcal{N}$ is defined to be the symmetrized sum
$E^{\rm sym}(\mathcal{M},\mathcal{N})=E^{\rm inj}(\mathcal{M},\mathcal{N})+E^{\rm inj}(\mathcal{N},\mathcal{M})\geq 0.$
Derksen, Weyman and Zelevinsky [DWZ10, Theorem 7.1] proved that the integer $E^{\rm sym}(\mathcal{M},\mathcal{N})$ is mutation-invariant in the sense that
$E^{\rm sym}(\mathcal{M},\mathcal{N})=E^{\rm sym}(\mu_{k}(\mathcal{M}),\mu_{k}(\mathcal{N})),$
where $\mu_{k}(\mathcal{M})$ and $\mu_{k}(\mathcal{N})$ are decorated representations of the new quiver with potential $\mu_{k}(Q,W)$ obtained by mutation in direction $k$. As a corollary, they obtained $E^{\rm sym}(\mathcal{M},\mathcal{M})=0$ for any negative-reachable decorated representations, i.e., those decorated representations corresponding to cluster monomials. This result plays a crucial role in their proofs of [DWZ10, Conjectures 1.3–1.6].
Let $(Q,W)$ be a quiver with potential. There is a dg algebra $\Gamma=\Gamma_{(Q,W)}$ and a triangulated category $\mathcal{C}_{(Q,W)}$ associated to $(Q,W)$, which are called (complete) Ginzburg dg algebra and cluster category, cf., [Amiot_2009] and [Keller-Yang_2011]. We remark that $\Gamma$ can be viewed as an object in $\mathcal{C}_{(Q,W)}$ satisfying $\operatorname{\mathsf{Hom}}_{\mathcal{C}_{(Q,W)}}(\Gamma,\Gamma[1])=0$. Plamondon [Plamondon_2011b, Section 4.1] constructed a bijection $\Psi_{(Q,W)}$ from the isomorphism classes of decorated representations of $(Q,W)$ to the isomorphism classes of objects in a full subcategory of $\mathcal{C}_{(Q,W)}$. For two decorated representations $\mathcal{M}$ and $\mathcal{N}$ of $(Q,W)$, we denote by $M=\Psi_{(Q,W)}(\mathcal{M})$ and $N=\Psi_{(Q,W)}(\mathcal{N})$ the corresponding objects in the cluster category $\mathcal{C}_{(Q,W)}$. In [Plamondon_2011b, Proposition 4.15], Plamondon proved
$E^{\rm inj}(\mathcal{M},\mathcal{N})=\dim(\Gamma[1])(M,N[1]),$
where $(\Gamma[1])(M,N[1])$ denotes the subspace of $\operatorname{\mathsf{Hom}}_{\mathcal{C}_{(Q,W)}}(M,N[1])$ consisting of morphisms that factor through an object of $\operatorname{\mathsf{add}}\nolimits(\Gamma[1])$. Thus we have
$E^{\rm sym}(\mathcal{M},\mathcal{N})=\dim(\Gamma[1])(M,N[1])+\dim(\Gamma[1])(N,M[1]).$
For the case of $\mathcal{M}=\mathcal{N}$, Plamondon also proved that the above equality becomes
$E^{\rm sym}(\mathcal{M},\mathcal{M})=\dim\operatorname{\mathsf{Hom}}_{\mathcal{C}_{(Q,W)}}(M,M[1]).$
One has that $E^{\rm sym}(\mathcal{M},\mathcal{M})=0$ if and only if the corresponding object $M$ in the cluster category $\mathcal{C}_{(Q,W)}$ is rigid, that is, $\operatorname{\mathsf{Hom}}_{\mathcal{C}_{(Q,W)}}(M,M[1])=0$. Notice that if $\mathcal{M}$ corresponds to a cluster monomial, then $M$ is rigid in $\mathcal{C}_{(Q,W)}$. Let us mention another very special case. If $N$ belongs to $\operatorname{\mathsf{add}}\nolimits(\Gamma)$, that is, $\mathcal{N}$ corresponds to a cluster monomial in initial cluster variables, then it is easy to see
(1.1) $E^{\rm sym}(\mathcal{M},\mathcal{N})=\dim\operatorname{\mathsf{Hom}}_{\mathcal{C}_{(Q,W)}}(M,N[1]).$
Inspired by Derksen-Weyman-Zelevinsky’s $E$-invariant for Jacobian algebras, the $E$-invariant for any finite dimensional algebra is introduced by Derksen and Fei [DK-2015] in the study of general presentations and by Adachi, Iyama and Reiten [air_2014] in the study of $\tau$-tilting theory.
Recently, tropical $F$-polynomials were studied by Fei [fei_2019a, fei_2019b] from the viewpoint of representation theory. It turns out that $E^{\rm inj}(\mathcal{M},\mathcal{N})$ is related to the evaluations of tropical $F$-polynomials, c.f., [fei_2019b, Theorem 3.22].
@@
$\mathfrak{d}$-invariant. The $\mathfrak{d}$-invariant is introduced by Kang, Kashiwara, Kim and Oh [kkko-2018] in the monoidal categorification of (quantum) cluster algebras using representations of quiver Hecke algebras. Here we will mainly follow [kkko-2018, kk_2019]. Let
[…]
$\Lambda(M,N)=-c.$
The $\mathfrak{d}$-invariant $\mathfrak{d}(M,N)$ associated to $(M,N)$ is defined to be
@@
[…]
@@
Recently, Kashiwara, Kim, Oh and Park [kkop-2020] introduced the similar invariants $\Lambda(M,N)$ and $\mathfrak{d}(M,N)$ into the monoidal categorification of cluster algebras using representations of quantum affine algebras. In particular, in this setting, they proved that (i) $\mathfrak{d}(M,N)$ takes non-negative values; (ii) if $M$ is a real simple object, then $\mathfrak{d}(M,M)=0$, cf. [kkop-2020, Proposition 3.16, Corollary 3.17].
Tropical invariant and $F$-invariant. In this paper, we introduce the $F$-invariant in cluster algebras using tropicalization. This is an analog of the $E$-invariant introduced by Derksen, Weyman and Zelevinsky in the additive categorification of cluster algebras and the $\mathfrak{d}$-invariant introduced by Kang, Kashiwara, Kim and Oh in the monoidal categorification of (quantum) cluster algebras. The $F$-invariant is also a great improvement of Fu-Gyoda’s $f$-compatibility degree defined on the set of cluster variables.
[…]
For any choice of seed $({\bf x}_{w},\widetilde{B}_{w})$, we define an integer $\langle u,u^{\prime}\rangle_{w}$ for a pair of good elements $u$ and $u^{\prime}$ in Definition 4.7. Our first main result is as follows:

### v4

@@ Abstract.
We consider skew-symmetrizable (upper) cluster algebras with a compatible Poisson structure, called $\mathsf{\Lambda}$-(upper) cluster algebras. For any two good elements (e.g., cluster monomials) in a $\mathsf{\Lambda}$-upper cluster algebra, we introduce two invariants, called tropical invariant and $F$-invariant. We prove that (i) the product of two cluster monomials is still a cluster monomial if and only if their $F$-invariant is zero; (ii) if two cluster variables are log-canonical, then they are contained in the same cluster; and (iii) the notion of $F$-invariant for a pair of cluster monomials can be defined for any (upper) cluster algebra, regardless of whether it is a $\mathsf{\Lambda}$-(upper) cluster algebra.
When restricting to cluster monomials, we prove that the tropical invariant and $F$-invariant respectively coincide with the $\Lambda$-invariant and twice $\mathfrak{d}$-invariant in the monoidal cluster categorification using various monoidal subcategories of finite-dimensional modules over quantum affine algebras and quiver Hecke algebras; and we prove that the $F$-invariant coincides with the $E$-invariant in the additive cluster categorification using the theory of quivers with potentials.
Inspired by $F$-invariant, we introduce the dominant sets for seeds of cluster algebras as a replacement of torsion classes for $\tau$-tilting pairs in $\tau$-tilting theory. With the help of the dominant sets, we prove that the oriented exchange graphs of cluster algebras are acyclic. In particular, this implies that green mutations induce a partial order on the set of seeds (up to seed equivalence) of cluster algebras. We prove that the oriented exchange graphs of cluster algebras coincide with the Hasse quivers of the above posets of seeds.
@@ 1. Introduction
1.1. Backgrounds
Cluster algebras were introduced by Fomin and Zelevinsky [fz_2002] as a combinatorial approach to the dual canonical bases (or upper global bases) of quantum groups and to the theory of total positivity in algebraic groups. Such algebras often arise as the coordinate rings of spaces, such as double Bruhat cells [bfz_2005], unipotent cells [gls_2011], double Bott-Samelson cells [Shen-Weng-2021] and arise as the Grothendieck rings of certain monoidal subcategories of the representations of quantum affine algebras [HL_2010, HL_2013, kkop-2024] and quiver Hecke algebras [kkko-2018].
@@ 1.1. Backgrounds
Compatibility degrees in cluster algebras
In the study of Y-systems and generalized associahedra, Fomin and Zelevinsky [fz-2003y] introduce a function
$(-\mid\mid-)\colon\Phi_{\geq-1}\times\Phi_{\geq-1}\rightarrow\mathbb{Z}_{\geq 0}$
defined on the set $\Phi_{\geq-1}$ of almost positive roots associated to a Cartan matrix $C$ of finite type. Let $\mathcal{A}$ be a cluster algebra of finite type with the same Cartan-Killing type as the Cartan matrix $C$. Fomin and Zelevinsky [fz_2003f] proved that there is a bijection from $\Phi_{\geq-1}$ to the set $\mathcal{X}$ of (unfrozen) cluster variables of $\mathcal{A}$ and they proved that two cluster variables $x_{\alpha}$ and $x_{\beta}$ are contained in the same cluster if and only if their corresponding roots $\alpha,\beta\in\Phi_{\geq-1}$ are compatible, i.e., $(\alpha\mid\mid\beta)=0$.
Now let $\mathcal{A}$ be any skew-symmetrizable cluster algebra (not necessarily finite type). In [cao-li-2020], Li and the author of the present paper proved that there exists a function
$(-\mid\mid-)_{d}\colon\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{Z}_{\geq-1}$
defined on the set of (unfrozen) cluster variables $\mathcal{X}$ of $\mathcal{A}$, which is called the $d$-compatibility degree. The values of the $d$-compatibility degree $(-\mid\mid-)_{d}$ are given by the components of the $d$-vectors [fomin_zelevinsky_2007] in cluster algebras. It is proved that two cluster variables $x,z\in\mathcal{X}$ are contained in the same cluster if and only if
@@
$(-\mid\mid-)_{f}\colon\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{Z}_{\geq 0}$
@@
defined on the set $\mathcal{X}$ of (unfrozen) cluster variables of $\mathcal{A}$. Based on the results in [cao-li-2020], Fu and Gyoda proved that two cluster variables $x,z\in\mathcal{X}$ are contained in the same cluster if and only if $(x\mid\mid z)_{f}=0$. When restricted to cluster algebras of finite type, the $f$-compatibility degree coincides with Fomin-Zelevinsky’s compatibility degree.
$\Lambda$-invariant and $\mathfrak{d}$-invariant
@@
The $\Lambda$-invariant and $\mathfrak{d}$-invariant are introduced by Kang, Kashiwara, Kim and Oh [kkko-2018] in the representation theory of quiver Hecke algebras and by Kashiwara, Kim, Oh and Park [kkop-2020] in the representation theory of quantum affine algebras in their study of the monoidal categorification of (quantum) cluster algebras.
Here we briefly recall these notions in the quiver Hecke algebras case and we mainly follow [kkko-2018, kk_2019]. Let
$R\text{-gmod}=\bigoplus_{\beta\in Q^{+}}R(\beta)\text{-gmod}$
be the category of finite-dimensional graded modules with homomorphisms of degree 0 given in [kk_2019, Page 2270], which is a monoidal abelian category with the tensor product $\circ$} given there. Denote by $q$ the grading shift functor in $R\text{-gmod}$: for $M\in R\text{-gmod}$, we have $(qM)_{i}=M_{i-1}$ for $i\in\mathbb{Z}$. For each pair $(M,N)$ with $0\neq M\in R(\beta)\text{-gmod}$ and $0\neq N\in R(\gamma)\text{-gmod}$, Kang, Kashiwara, Kim and Oh [kkko-2018, Definition 2.2.1] constructed a renormalized $r$-matrix ${\bf r}_{M,N}$, which is a morphism from $M{\circ}N$ to $q^{c}N{\circ}M$ in $R\text{-gmod}$ for some integer $c$. The $\Lambda$-invariant $\Lambda(M,N)$ of the pair $(M,N)$ is defined as the homogeneous degree of ${\bf r}_{M,N}$, i.e.,
$\Lambda(M,N)=-c\;\in\mathbb{Z}.$
The $\mathfrak{d}$-invariant $\mathfrak{d}(M,N)$ of the pair $(M,N)$ is defined as half of the symmetrized sum:
@@
It turns out that $\mathfrak{d}(M,N)\in\mathbb{Z}_{\geq 0}$, cf. [kkko-2018, Lemma 3.2.1].
Remark 1.1.
Denote by $R\text{-mod}^{0}$ the category obtained from $R\text{-gmod}$ by forgetting the gradings. By construction of the renormalized $r$-matrices, we have
$\Lambda(q^{a}M,q^{b}N)=\Lambda(M,N)\;\;\;\text{and}\;\;\;\mathfrak{d}(q^{a}M,q^{b}N)=\mathfrak{d}(M,N),\;\;\forall a,b\in\mathbb{Z}.$
Thus the $\Lambda$-invariant and $\mathfrak{d}$-invariant descend to the ungraded category $R\text{-mod}^{0}$.
A simple object $S$ in a monoidal abelian category is said to be real, if the tensor product of $S$ with itself is still simple. Kang, Kashiwara, Kim and Oh proved that if $M$ is a real simple object in $R\text{-gmod}$, then $\mathfrak{d}(M,M)=0$, cf. [kkko-2018, Lemma 3.2.3]. In monoidal categorification of cluster algebras, each cluster monomial corresponds to a real simple object. In particular, the $\mathfrak{d}$-invariant associated to a “cluster monomial” is zero.
@@
Now let us turn to the $\mathfrak{d}$-invariant in the representation theory of quantum affine algebras. In this setting, Kashiwara, Kim, Oh and Park [kkop-2020] proved that (i) $\mathfrak{d}(M,N)$ takes non-negative values; (ii) if $M$ is a real simple object, then $\mathfrak{d}(M,M)=0$, cf. [kkop-2020, Proposition 3.16, Corollary 3.17].
$E$-invariant
The $E$-invariant is introduced by Derksen, Weyman and Zelevinsky [DWZ10] in the additive categorification of cluster algebras using decorated representations of quivers with potentials. For any two decorated representations $\mathcal{M}$ and $\mathcal{N}$ of a quiver with potential $(Q,W)$, Derksen, Weyman and Zelevinsky [DWZ10, Section 7] defined an integer $E^{\rm inj}(\mathcal{M},\mathcal{N})$, which is called the partial $E$-invariant of the pair $(\mathcal{M},\mathcal{N})$ in this paper (see Definition 6.2). The $E$-invariant $E^{\rm sym}(\mathcal{M},\mathcal{N})$ of the pair $(\mathcal{M},\mathcal{N})$ is defined to be the symmetrized sum:
$E^{\rm sym}(\mathcal{M},\mathcal{N})=E^{\rm inj}(\mathcal{M},\mathcal{N})+E^{\rm inj}(\mathcal{N},\mathcal{M})\;\;\in\mathbb{Z}.$
It is known from [DWZ10, Section 10] (or [CLS-2015, Section 3]) that
$E^{\rm inj}(\mathcal{M},\mathcal{N})\geq 0\;\;\;\text{and}\;\;\;E^{\rm sym}(\mathcal{M},\mathcal{N})\geq 0.$
Derksen, Weyman and Zelevinsky [DWZ10, Corollary 7.2] proved that $E^{\rm sym}(\mathcal{M},\mathcal{M})=0$ for any negative-reachable decorated representations, i.e., those decorated representations corresponding to cluster monomials.
Inspired by Derksen-Weyman-Zelevinsky’s $E$-invariant for quivers with potentials, i.e., for Jacobian algebras, the $E$-invariant for any finite-dimensional algebra is introduced by Derksen and Fei [DK-2015] in the study of general presentations and by Adachi, Iyama and Reiten [air_2014] in the study of $\tau$-tilting theory.
1.2. Results (I): Tropical invariant and $F$-invariant
We consider skew-symmetrizable (upper) cluster algebras with a compatible Poisson structure, called $\mathsf{\Lambda}$-(upper) cluster algebras. In a $\mathsf{\Lambda}$-upper cluster algebra, each seed $({\bf x}_{t},(\widetilde{B}_{t})_{m\times n})$ is endowed with a skew-symmetric integer matrix $(\Lambda_{t})_{m\times m}$, called Poisson coefficient matrix such that
(1.1) $\widetilde{B}_{t}^{T}\Lambda_{t}=(S\mid{\bf 0})$
for some fixed diagonal matrix $S=diag(s_{1},\ldots,s_{n})$ with $s_{i}>0$. We call the triple $({\bf x}_{t},\widetilde{B}_{t},\Lambda_{t})$ a $\mathsf{\Lambda}$-seed of $\mathcal{U}$.
For any two good elements (e.g., cluster monomials) in a $\mathsf{\Lambda}$-upper cluster algebra, we introduce two invariants, called tropical invariant and $F$-invariant, which generalize the $\Lambda$-invariant and $\mathfrak{d}$-invariant in the monoidal cluster categorification. The $F$-invariant also generalizes the $E$-invariant in the additive cluster categorification.
A good element is a compatibly pointed element in the sense of [qin_2019] satisfying the universally positive property (see Definition 3.6). We view such elements as a replacement of decorated representations of quivers with potentials and finite-dimensional graded modules of quiver Hecke algebras.
The Laurent expansion of a good element $u\in\mathcal{U}$ with respect to any seed $({\bf x}_{w},\widetilde{B}_{w})$ of $\mathcal{U}$ has a canonical expression:
$u={\bf x}_{w}^{{\bf g}_{u}^{w}}F_{u}^{w}(\hat{y}_{1;w},\ldots,\hat{y}_{n;w})$
where ${\bf g}_{u}^{w}\in\mathbb{Z}^{m}$ and $F_{u}^{w}\in\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$ is a polynomial with constant term $1$. They are respectively called the extended $g$-vector and $F$-polynomial of $u$ with respect to seed $({\bf x}_{w},\widetilde{B}_{w})$.
For each good element $u$ in a $\mathsf{\Lambda}$-upper cluster algebra $\mathcal{U}$, we introduce a semifield homomorphism
$\beta_{u}\colon\mathbb{F}_{>0}\coloneqq\mathbb{Q}_{\rm sf}(x_{1;t_{0}},\ldots,x_{m;t_{0}})\rightarrow\mathbb{Z}^{\rm max},$
where $\mathbb{F}_{>0}$ is the universal semifield generated by the initial cluster variables of $\mathcal{U}$ and $\mathbb{Z}^{\max}=(\mathbb{Z},+,\max)$ is the tropical semifield (see Proposition 4.4). Using such semifield homomorphisms, we introduce the tropical invariant and $F$-invariant.
Definition 1.2 (Tropical invariant and $F$-invariant).
Let $u$ and $u^{\prime}$ be two good elements in a $\mathsf{\Lambda}$-upper cluster algebra $\mathcal{U}$. The tropical invariant $\langle u,u^{\prime}\rangle_{\rm trop}$ and $F$-invariant $(u\mid\mid u^{\prime})_{F}$ of the pair $(u,u^{\prime})$ are defined as follows:
$\langle u,u^{\prime}\rangle_{\rm trop}\coloneqq\beta_{u^{\prime}}(u)\;\;\;\text{and}\;\;\;(u\mid\mid u^{\prime})_{F}=\beta_{u^{\prime}}(u)+\beta_{u}(u^{\prime})=\langle u,u^{\prime}\rangle_{\rm trop}+\langle u^{\prime},u\rangle_{\rm trop}.$
Given a non-zero polynomial
and a vector ${\bf r}\in\mathbb{Z}^{n}$, we denote by
$F[{\bf r}]\coloneqq\max\{{\bf v}^{T}{\bf r}\mid c_{{\bf v}}\neq 0\}\;\in\mathbb{Z}.$
It is easy to see if $F$ has the constant term $1$, then $F[{\bf r}]\in\mathbb{Z}_{\geq 0}$ for any ${\bf r}\in\mathbb{Z}^{n}$.
The first main result in this paper is that for each $\mathsf{\Lambda}$-seed $({\bf x}_{w},\widetilde{B}_{w},\Lambda_{w})$ of $\mathcal{U}$, we have an explicit formula to calculate the tropical invariant and $F$-invariant.